# Domain Adaptation-Enhanced Searchlight: Enabling brain decoding from visual perception to mental imagery

### v1

@@ 6.1 Comparison of DA techniques
Figure 1 shows the unique Critical Difference (CD) diagram obtained by concatenating all the observations (both across $N_{t}$ and across different subjects). The scale above shows the average ranking of each algorithm across all observations, and the horizontal lines group together algorithms that are not statistically different. From this, we conclude that RTLC was the best algorithm, followed by the group of NAIVE, TAB, BW and FA which are statistically indistinguishable ($\alpha=0.05$). On the other hand, SA is shown to be the worst.
@@
Figure 2 summarizes the results of the tests in the 18 CD diagrams arising from concatenating the 10 tables corresponding to different $N_{t}$ values for each subject, which are shown in Supplementary Figure 9. For each cell $(i,j)$ in this table, we report the number of subjects for which algorithm $j$ was significantly better than algorithm $i$ ($\alpha=0.05$). This can be extracted from the CD diagrams by counting how many times $j$ appears to the left of $i$ and substracting the number of times they are joined by a horizontal line. The sum of each column represents the number of times each algorithm was better than others and, complementarily, the sum of each row reports the number of times each algorithm was significantly worse than others.
Figure 2: Frequency table. For each cell $(i,j)$, the numerical value shows for how many subjects algorithm $j$ was significantly superior to algorithm $i$. The sum of each column accounts for the number of times each algorithm was significantly better than others, and the sum of each row shows how many times each algorithm was significantly worse than others.
@@
  • 1.
    RTLC outperformed others a total of 127 times, and was outperformed 51 times. It was the most outperforming and least outperformed algorithm.
  • 2.
@@
  • 4.
    NAIVE was better than BASELINE for 10 subjects and worse for 8.
  • 5.
    SA was the most outperformed and least outperforming algorithm.
@@ 6.3 DA-enhanced searchlight analysis
Figure 6 shows the projection of the average balanced accuracy of each approach (computed across the subjects) onto the areas of statistically significant brain decoding. We see that only within-domain prediction attains good balanced accuracy when averaged across all the subjects, and that in this case (perception) decoding is possible in a highly distributed set of brain regions. For the baseline and naïve approaches, decoding is only possible in a small isolated region. On the contrary, RLTC is able to obtain significantly above chance imagery decoding in an extensively spread brain volume.
Figure 7 shows the results of the statistical comparison between RTLC and the naïve approach, highlighting only the statistically significant clusters ($p<0.05$, whole-brain corrected). The naïve method was never statistically significantly above the baseline.
Figure 8 accounts for the inter-subject variability of RTLC performance. Each set of brain slices corresponds to a different subject, and shows the average accuracy across the 100 data partitions projected onto the space where such average is statistically greater than chance level. Supplementary Figures 10, 11 and 12, respectively, show the subject variability of the perception decoding, the baseline cross-domain decoding and the cross-domain decoding with naïve inclusion of target instances in the training set.
Figure 6: Between-subject average of the balanced accuracy of the different approaches in the regions that obtained statistically above chance decoding.
@@
Figure 8: Inter-subject variability of the DA-enhaced searchlight accuracy. For each subject, the color of each voxel represents the average RTLC accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically above 0.5.
@@ 7 Discussion
We hypothesized that predictive models trained with fMRI data from different experimental conditions could benefit from domain adaptation (DA). Our results support this hypothesis, showing that DA using the RTLC technique significantly improves cross-domain brain decoding from perception to imagery compared to the naive inclusion of imagery instances in the training set. Furthermore, 6 out of the 11 DA methods studied outperformed the baseline (see Figure 1). With our DA-enhanced searchlight approach, we demonstrated statistically significant pattern similarities between perception and imagery in the visual cortex and beyond, including a distributed set of brain regions in the frontoparietal cortex.
@@
Our results show that RTLC can be beneficial for cross-domain brain decoding when employed in the union of the ROIs under study, and also when integrated into a searchlight procedure. These results align with previous work stating that BCIs can be improved by incorporating DA into their training methodology [64, 65]. This requires little extra effort, aside from gathering a small sample of data from the target domain. In the case of fMRI studies, acquiring this data involves no significant additional cost. Importantly, we have shown that naïvely concatenating this data into the training set of a conventional classifier is less effective than using RTLC to incorporate the same information.
@@
In the analysis across individual ROIs presented in Section 6.2, we assess the contribution of the different ROIs to RTLC performance, as well as their informativeness according to the baseline. We see that RTLC performs best in the FFG, but is also able to extract information from ROIs that are indirectly related to vision such as the inferior parietal lobe (IPL), the inferior temporal gyrus (ITG), and the middle temporal gyrus (MTG).
@@
The success of DA is strongly limited by the signal-to-noise ratio of the data. When the task of interest is difficult in the source domain itself, resulting in unreliable starting models, generalization to the target domain will be poor regardless of the DA technique employed. In our case, not all ROIs contain valuable information on visual object category, which becomes evident in Figure 4 and in the searchlight analysis. Specifically, Figure 4 highlights the fusiform gyrus (FFG) as the most informative ROI for visual perception decoding in all cases, whereas ROIs such as the inferior frontal gyrus, pars orbitalis (IFGorbital), the frontopolar cortex (FP) and the superior frontal gyrus (SFG), among others, encode little information about visual object category. Equivalently, from Supplementary Figure 10 we conclude that all subjects have good perception-to-perception decoding in the areas of the occipital lobe associated with vision, but only some subjects encode visual information in additional brain regions. For around half of the subjects, the visual information is distributed on the whole brain above chance level, even though the decoding is always stronger in the occipital lobe. Hence, the between subject average decoding is predominant in the occipital lobe but also extends to other areas (Figure 6 - perception).
@@
When it comes to the direct applicability of standard perception-based models to the imagery domain, Figure 6 (baseline and naïve) reveals a poor transferability on average. Inspecting Supplementary Figures 11 and 12, we discover that the extent of the domain shift (indicated by the performance gap of these approaches compared to perception prediction) is in fact very subject-dependent. Participants such as Subject 7 and Subject 14 exhibit good transferability, but others such as Subject 5 or Subject 8 have very sparse above-chance decoding regions in imagery. Crucially, we reiterate that the across-subject average of the decoding is limited to a very small brain region, which makes the standard models unreliable for additional out-of-sample subjects in the imagery domain.
@@
In contrast, even when averaged across the subjects, DA-enhanced models attain above chance decoding in mostly the same areas as perception prediction does, although, expectedly, the acuraccy drops considerably. Once again, this is explained by the subject variability depicted in Figure 8. Some subjects benefit from DA more than others do. Participants such as Subject 8 remain a challenge, whereas for others like Subject 5 the adapted models are well tailored to the imagery scenario highlighting imagery-specific brain regions. In cases such as Subject 5, Subject 13 and Subject 10 the parietal lobe appears to play no role on encoding visual information that distinguishes between the living and non-living classes (see Figure 10), but it contains relevant imagery information which is discriminative of those classes.
@@
Regarding why RTLC outperforms other DA methods in the ROI-based comparison, the simplicity of the technique might make it adequate to deal with high-dimensional data. In our comparison of DA techniques, the amount of examples is low compared to the amount of voxels (see Table 3 for the exact numbers). This may be detrimental to more sophisticated DA techniques, but still feasible for RTLC. In addition, RTLC is tailored to the specific estimator $h$ used for the source domain. Thus, if $h$ has good performance, it will be easier to obtain a good $h_{RTLC}$. This is evidenced in Supplementary Figure 10 and in Figure 8, which respectively indicate the performance of the searchlight $h$ and $h_{RTLC}$.
@@
In fact, from the point of view of ML, it is important to analyze the limitations of current DA techniques in real-word scenarios. Usually, DA methods are tested considering artificial benchmarks such as Office-31 [66] and MNIST [67] or real problems with a low dimensionality. fMRI data offers the opportunity to evaluate these methods, elucidating whether the complexity of the methods is translated into an enhanced performance. Therefore, one of the goals of this paper is to evaluate the performance of state of the art DA techniques when dealing with high-dimensional, noisy and complex data.
@@
Among the limitations of our work, we lack a comprehensive characterization of the participants, which was not possible to achieve with the available data. Our results reveal a large amount of subject variability both in terms of the performance of baseline classifiers and the success of DA efforts, which is natural given the individuality of brain processes, but which we are not able to adequately explain. For instance, measuring the subjective vividness of visual imagery as done in [68] would have allowed us to establish the relationship between individual imagery strategies and the degree of improvement brought by DA. Another limitation is that some DA methods that did not excel at the ROI-level comparison could have been even better than RTLC in the searchlight analysis, which does not suffer from the curse of dimensionality. However, due to the computationally intensive nature of the searchlight procedure, a comparison of all the DA techniques on a voxel level would have been unattainable. This work provides a novel precedent of fMRI DA-enhanced searchlight for cross-domain brain decoding, while acknowledging that RTLC is not necessarily the best performing DA technique in this context. Nevertheless, sophisticated DA techniques and/or underlying estimators such as neural networks are impractical to integrate into a searchlight procedure that requires fitting tens of thousands of such models for each data partition.
The idea of a DA-enhanced searchlight could also be applied to other domain-shift scenarios. Future research lines include testing the validity of this approach in cross-subject and cross-device problems, where the input space is not shared across domains. This additional difficulty could be addressed with proper representation learning techniques.
@@ 8 Conclusions
In this paper, we have introduced an effective way to integrate DA into the searchlight procedure, significantly enhancing classification performance in cross-domain brain decoding. Through a comprehensive analysis involving data from 14 ROIs, we demonstrated that several DA techniques surpassed our baseline model, which was trained exclusively on perception data and evaluated on imagery data. This underscores the presence of a domain shift between the neural activities associated with visual perception and imagery of living versus non-living objects.
The best performing method was RTLC. Utilizing RTLC, we assessed both the nature and location of the domain shift within our DA-enhanced searchlight approach. Remarkably, RTLC enabled above-chance decoding of imagined stimuli in a highly distributed set of brain regions, including the visual cortex and the frontoparietal cortex. This contrasts sharply with non-DA searchlight procedures, which did not achieve comparable performance.
@@
9 Declaration of Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
10 CRediT author contributions
Olza: Conceptualization, Investigation, Software, Writing - Original draft Soto: Conceptualization, Data curation, Writing - review and editing, Funding acquisition, Supervision Santana: Conceptualization, Writing - review and editing, Funding acquisition, Supervision
11 Acknowledgements
This work has been supported by the Basque Government via the IKUR strategy, project IT1504-22 and the BERC 2022-2025 program, by the Spanish Ministry of Science and Innovation (PID2022-137442NB-I00), by Elkartek (KK-2023/00090), by the Spanish Ministry of Economy and Competitiveness, through the ’Severo Ochoa’ Programme for Centres/Units of Excellence (CEX2020-001010-S) and also from project grant PID2019-105494GB-I00.

### v3

@@ 5.1 Comparison of DA techniques
Figure 2 shows the unique Critical Difference (CD) diagram obtained by concatenating all the observations (both across $N_{t}$ and across different subjects). The scale above shows the average ranking of each algorithm across all observations, and the horizontal lines group together algorithms that are not statistically different. From this, we conclude that RTLC was the best algorithm, followed by the group of NAIVE, TAB, BW and FA which are statistically indistinguishable ($\alpha=0.05$). On the other hand, FT is shown to be the worst.
@@
Figure 3 summarizes the results of the tests in the 18 CD diagrams arising from concatenating the 10 tables corresponding to different $N_{t}$ values for each subject, which are shown in Supplementary Figure 10. For each cell $(i,j)$ in this table, we report the number of subjects for which algorithm $i$ was significantly better than algorithm $j$ ($\alpha=0.05$). This can be extracted from the CD diagrams by counting how many times $i$ appears to the left of $j$ and subtracting the number of times they are joined by a horizontal line. The sum of each row represents the number of times each algorithm was better than others and, complementarily, the sum of each column reports the number of times each algorithm was significantly worse than others.
Figure 3: Frequency table. For each cell $(i,j)$, the numerical value shows for how many subjects algorithm $i$ was significantly superior to algorithm $j$. The sum of each row accounts for the number of times each algorithm was significantly better than others, and the sum of each column shows how many times each algorithm was significantly worse than others.
@@
  • RTLC outperformed others a total of 169 times, and was outperformed 60 times. It was the most outperforming and least outperformed algorithm.
@@
  • NAIVE was better than BASELINE for 10 subjects and worse for 7.
  • FT was the most outperformed and least outperforming algorithm.
@@ 5.2 Analysis of individual ROIs using RTLC
In the case of the “Generic Object Decoding” dataset, the best and worst methods were respectively PRED and DCORAL (Supplementary Figure 14). FA, BW and TAB were also better than NAIVE, although with no statistical significance. PRED outperformed other algorithms 47 times and was only outperformed once (Supplementary Figure 15). As detailed in B, the statistical power of the tests was diminished because of the lower number of subjects and feasible data partitions. Although no statistical significance was obtained, Supplementary Figure 16 reveals that FA and PRED were better than NAIVE in 4 subjects, BW for 3 and TAB for 2.
@@ 5.3 DA-enhanced searchlight analysis
Figure 7 shows the projection of the average balanced accuracy of each approach (computed across the subjects) onto the areas of statistically significant brain decoding. We see that only within-domain prediction attains good balanced accuracy when averaged across all the subjects, and that in this case (perception) decoding is possible in a highly distributed set of brain regions. For the baseline and naïve approaches, decoding is only possible in a small, isolated region. On the contrary, our DA-enhanced searchlight with either RLTC and BW is able to obtain significantly above chance imagery decoding in an extensively spread brain volume.
Figure 8 shows the results of the statistical comparison between RTLC and the naïve approach using a two-sample permutation test, highlighting only the statistically significant clusters ($p<0.05$, whole-brain corrected). The naïve method was never statistically significantly above the baseline. Supplementary Figure 17 compares BW and naïve. Remarkably, a two-sample permutation test reveals that all the significant clusters that were found with RTLC are still present with BW, highlighting the robustness of the neuroscientific conclusions.
Figure 9 accounts for the inter-subject variability of RTLC performance. Conversely, Supplementary Figure 19 displays BW performance for each subject. Each set of brain slices corresponds to a different subject, and shows the average accuracy across the 100 data partitions projected onto the space where such average is statistically greater than chance level. Supplementary Figures 11, 12 and 13, respectively, show the subject variability of the perception decoding, the baseline cross-domain decoding and the cross-domain decoding with naïve inclusion of target instances in the training set.
As illustrated in Supplementary Figure 18, the cross-subject average accuracy achieved using BW significantly exceeded chance levels across an even broader range of brain regions compared to RTLC. Notably, with BW, our DA-enhanced searchlight successfully classifies imagery of Living versus Non-Living items in regions that extend into the frontal lobe. However, in terms of symmetry between the left and right hemispheres, the accuracy map for BW exhibits less balance compared to RTLC. This asymmetry is likely attributable to cross-subject averaging, as indicated by the comparison between Figure 9 and Supplementary Figure 19. Specifically, BW demonstrates poor decoding performance for subjects 9, 15, and 17, whereas RTLC achieves successful decoding for these individuals.
Figure 7: Between-subject average of the balanced accuracy of the different approaches in the regions that obtained statistically above chance decoding with RTLC.
@@
Figure 9: Inter-subject variability of the DA-enhanced searchlight accuracy. For each subject, the colour of each voxel represents the average RTLC accuracy over the 100 repetitions of the experiment in the regions where such accuracy is statistically above 0.5.
@@ 6 Discussion
We hypothesized that predictive models trained with fMRI data from different experimental conditions could benefit from DA. Our results support this hypothesis, showing that DA significantly improves cross-domain brain decoding from perception to imagery compared to the naive inclusion of imagery instances in the training set. On an extensive ROI level, statistically significant improvement was attained for RTLC in the binary dataset and for PRED in the multiclass setting. Furthermore, 6 out of the 16 DA methods studied significantly outperformed the baseline in our dataset (see Figure 2), and 8 methods significantly outperformed the baseline in the multiclass dataset (Supplementary Figure 14).
For both datasets, BW and TAB were statistically equivalent to NAIVE in the ROI-based comparison, suggesting that these methods may perform even better at a local level, where the curse of dimensionality is absent. Indeed, the DA-enhanced searchlight analysis using BW reveals the potential of this approach compared to NAIVE. Conversely, TAB is less suitable for DA-enhanced searchlight analysis due to its iterative nature and high computational cost. Given the unique characteristics of each dataset, we emphasize the importance of comparing multiple DA methods to guide the selection of the most effective technique for DA-enhanced searchlight, while considering both computational resources and the possibility that DA methods may perform optimally at a local level rather than at the ROI level.
With our DA-enhanced searchlight approach, we demonstrated statistically significant pattern similarities between perception and imagery in the visual cortex and beyond, including a distributed set of brain regions in the frontoparietal cortex.
@@
Our results show that DA can be beneficial for cross-domain brain decoding when employed in the union of the ROIs under study, and also when integrated into a searchlight procedure. These results align with previous work stating that BCIs can be improved by incorporating DA into their training methodology [71, 72]. This requires little extra effort, aside from gathering a small sample of data from the target domain. In the case of fMRI studies, acquiring this data involves no significant additional cost. Importantly, we have shown that naïvely concatenating this data into the training set of a conventional classifier is less effective than using DA methods to incorporate the same information.
@@
As mentioned above, the best performing DA method for our dataset was RTLC. In the analysis across individual ROIs presented in Section 5.2, we assess the contribution of the different ROIs to RTLC performance, as well as their informativeness according to the baseline. We see that RTLC performs best in the FFG, but is also able to extract information from ROIs that are indirectly related to vision such as the inferior parietal lobe (IPL), the inferior temporal gyrus (ITG), and the middle temporal gyrus (MTG).
@@
The success of DA is strongly limited by the signal-to-noise ratio of the data. When the task of interest is difficult in the source domain itself, resulting in unreliable starting models, generalization to the target domain will be poor regardless of the DA technique employed. In our case, not all ROIs contain valuable information on visual object category, which becomes evident in Figure 5 and in the searchlight analysis. Specifically, Figure 5 highlights the fusiform gyrus (FFG) as the most informative ROI for visual perception decoding in all cases, whereas ROIs such as the inferior frontal gyrus, pars orbitalis (IFGorbital), the frontopolar cortex (FP) and the superior frontal gyrus (SFG), among others, encode little information about visual object category. Equivalently, from Supplementary Figure 11 we conclude that all subjects have good perception-to-perception decoding in the areas of the occipital lobe associated with vision, but only some subjects encode visual information in additional brain regions. For around half of the subjects, the visual information is distributed on the whole brain above chance level, even though the decoding is always stronger in the occipital lobe. Hence, the between-subject average decoding is predominant in the occipital lobe but also extends to other areas (Figure 7 - perception).
@@
When it comes to the direct applicability of standard perception-based models to the imagery domain, Figure 7 (baseline and naïve) reveals a poor transferability on average. Inspecting Supplementary Figures 12 and 13, we discover that the extent of the domain shift (indicated by the performance gap of these approaches compared to perception prediction) is in fact very subject-dependent. Participants such as Subject 7 and Subject 14 exhibit good transferability, but others such as Subject 5 have sparse above-chance decoding regions in imagery. Crucially, we reiterate that the across-subject average of the decoding is limited to a very small brain region, which makes the standard models unreliable for additional out-of-sample subjects in the imagery domain.
@@
In contrast, even when averaged across the subjects, DA-enhanced models attain above chance decoding in mostly the same areas as perception prediction does, although, expectedly, the accuracy drops considerably. Once again, this is explained by the subject variability depicted in Figure 9 and Supplementary Figure 19. Some subjects benefit from DA more than others do. Participants such as Subject 17 remain a challenge, whereas for others like Subject 5 the adapted models are well tailored to the imagery scenario, highlighting imagery-specific brain regions. In cases such as Subject 5, Subject 13 and Subject 15 the parietal lobe appears to play no role on encoding visual information that distinguishes between the living and non-living classes (see Figure 11), but it contains relevant imagery information which is discriminative of those classes.
@@
Regarding why RTLC and PRED outperform other DA methods in the binary and multiclass datasets respectively in the ROI-based comparison, the simplicity of these techniques might make them adequate to deal with high-dimensional data. In both datasets, our comparison of DA techniques suffers from the curse of dimensionality. The amount of examples is low compared to the amount of voxels (see Table 3 for the exact numbers in our dataset). This may be detrimental to more sophisticated DA techniques, but still feasible for RTLC and PRED. In addition, both RTLC and PRED are tailored to the specific estimator $h$ used for the source domain. Explicitly, RTLC works as a regularized maximum likelihood estimator in the target domain penalized by the distance of the parameters to the optimal source domain parameters, and PRED works by augmenting the features of the source estimator $h$ with its own target domain predictions and then fitting the augmented estimator to the target domain. Thus, if $h$ has good performance, it will be easier to obtain a good $h_{RTLC}$ or $h_{PRED}$. For our dataset, this is evidenced in Supplementary Figure 11 and in Figure 9, which respectively indicate the performance of the searchlight $h$ and $h_{RTLC}$.
The BW technique also performed remarkably in the DA-enhanced searchlight, despite not attaining a statistically significant improvement over NAIVE at the ROI level. Once again, there are some similarities between the methodology of RTLC and BW. BW works by fitting $h_{D}A$ directly on source and target labelled data according to the modified loss $(1-\gamma)\mathcal{L}(h(X_{s}),y_{s})+\gamma\mathcal{L}(h(X_{t}),y_{t})$, where the parameter $\gamma$ controls the trade-off between fitting the source or the target data. In a way, this is related to RTLC because it minimizes a combined source-target loss where the target component can also be considered a penalty term, although RTLC does it in two consecutive steps.
@@
In fact, from the point of view of ML, it is important to analyse the limitations of current DA techniques in real-world scenarios. Usually, DA methods are tested considering artificial benchmarks such as Office-31 [73] and MNIST [74] or real problems with a low dimensionality. fMRI data offers the opportunity to evaluate these methods, elucidating whether the complexity of the methods is translated into an enhanced performance. Therefore, one of the goals of this paper is to evaluate the performance of state-of-the-art DA techniques when dealing with high-dimensional, noisy and complex data.
@@
Among the limitations of our work, we lack a comprehensive characterization of the 18 participants of our dataset, which was not possible to achieve with the available data. Our results reveal a large amount of subject variability both in terms of the performance of baseline classifiers and the success of DA efforts, which is natural given the individuality of brain processes, but which we are not able to adequately explain. For instance, measuring the subjective vividness of visual imagery as done in [75] would have allowed us to establish the relationship between individual imagery strategies and the degree of improvement brought by DA. Another limitation is that some DA methods that did not excel at the ROI-level comparison could have been even better than RTLC in the searchlight analysis, which does not suffer from the curse of dimensionality. In fact, this has been the case for BW, which performed even better than RTLC in the DA-enhanced searchlight. However, due to the computationally intensive nature of the searchlight procedure, a comparison of all the DA techniques on a voxel level would have been unattainable. This work provides a novel precedent of fMRI DA-enhanced searchlight for cross-domain brain decoding, while acknowledging that RTLC is not necessarily the best performing DA technique in this context. Nevertheless, sophisticated DA techniques and/or underlying estimators such as neural networks are impractical to integrate into a searchlight procedure that requires fitting tens of thousands of such models for each data partition. Another limitation is the lack of suitable alternative datasets to strengthen the neuroscientific implications of our findings. We were able to demonstrate the benefits of DA for visual perception to imagery transferability in the visual cortex ROI in a publicly available multiclass dataset, but this dataset is not appropriate for DA-enhanced searchlight due to various factors. To begin with, such dataset contains only 5 subjects, which is a small sample size for the permutation tests that we used to quantify and statistically validate the effect of DA across subjects. More importantly, transforming the searchlight results of all subjects to a common space (such as the MNI) is paramount to draw any conclusions, and the necessary transformation matrices to achieve this are not available with the “Generic Object Decoding” dataset. From the perspective of cognitive neuroscience, this paper is focused on investigating commonalities between visual perception and imagery, and the available datasets on this matter are scarce. However, the methodology of the DA-enhanced searchlight introduced in this work can be applied to any domain shift problem in the realm of neuroscience, opening the doors for future research.
For instance, inter-individual differences in brain structure and functional organization mean that a model trained on fMRI data from one person may not generalize well to others without substantial adjustment. These inter-subject particularities may be more prominent in subjective tasks such as imagery or memory. These cross-subject domain-shift scenarios represent a critical problem in which DA may prove useful to improve the generalizability and robustness of models trained to decode brain activity patterns in standard anatomical spaces. Handling cross-subject differences often involves transforming data into a common anatomical or functional space (in our case, the MNI space), but even with alignment, inter-subject variability remains a major factor limiting generalization in fMRI studies. Although cross-subject DA is beyond the scope of our paper, extending the DA-enhanced searchlight procedure to this setting would be important in future studies.
In fact, our results highlight the subject-dependency of imagery processes, where inter-subject heterogeneity is much more acute than in visual stimuli processing (as proven by comparing Supplementary Figure A.11 against Supplementary Figure A.12). By applying our methodology on each subject independently, we have managed to gain knowledge about each participant’s characteristics (Supplementary Figure 9 and Supplementary Figure A.10). Additionally, by transforming the results to MNI space and running permutation-based statistical tests, we have obtained general insights on the brain functioning of our study population.
@@ 7 Conclusions
In this paper, we have introduced an effective way to integrate DA into the searchlight procedure, significantly enhancing classification performance in cross-domain brain decoding.
Through a comprehensive analysis involving data from 14 ROIs and 18 subjects for a classification task distinguishing between living and non-living objects, we demonstrated that several DA techniques outperformed our baseline model. The baseline model was trained solely on perception data and tested on imagery data. Similar results were observed on a publicly available multiclass dataset, highlighting a clear domain shift between neural activities associated with visual perception and imagery.
The best performing method on our dataset was RTLC. Utilizing RTLC, we assessed both the nature and location of the domain shift within our DA-enhanced searchlight approach. Remarkably, RTLC enabled above-chance decoding of imagined stimuli in a highly distributed set of brain regions, including the visual cortex and the frontoparietal cortex. This contrasts sharply with non-DA searchlight procedures, which did not achieve comparable performance. To strengthen the robustness of our neuroscientific findings, we replicated the results using a DA-enhanced searchlight approach with an alternative DA technique (BW).
@@
Declarations
  • Availability of data and materials: The code for this paper is available at the GitHub repository AlexOlza/DA-enhanced-searchlight. The binary perception/imagery dataset analysed during the current study is available in the Open Science Framework (OSF) repository, https://osf.io/audmx/. The “Generic Object Decoding” dataset that supports the findings of this study is available in Figshare with the identifier https://figshare.com/articles/dataset/Generic_Object_Decoding/7387130.
  • Competing interests: The authors declare that they have no competing interests.
  • Funding: This work has been supported by the Basque Government via the IKUR strategy, project IT1504-22 and the BERC 2022-2025 program, by the Spanish Ministry of Science and Innovation (PID2022-137442NB-I00), by Elkartek (KK-2023/00090), by the Spanish Ministry of Economy and Competitiveness, through the ’Severo Ochoa’ Programme for Centres/Units of Excellence (CEX2020-001010-S) and also from project grant PID2019-105494GB-I00.
  • Author’s contributions: AO participated in the conceptualization and research of the methodology, implemented the software and wrote the original draft. DS was involved in the conceptualization, data curation, funding acquisition, supervision and revision of the manuscript. RS contributed to the conceptualization, funding acquisition, supervision and revision of the manuscript. All authors read and approved the final manuscript.
  • Acknowledgements: We are grateful to the anonymous subjects that took part in experimental data acquisition.
8 List of abbreviations
  • DA: Domain Adaptation
  • MRI: Magnetic Resonance Imaging
  • fMRI: Functional MRI
  • MVPA: Multivoxel (or multivariate) Pattern Classification Analysis
  • BCI: Brain-Computer Interface
  • ROI: Region of Interest
  • ML: Machine Learning
  • EEG: Electroencephalography
  • CNN: Convolutional Neural Network
  • MNI: Montreal Neurological Institute
  • LR: Logistic Regression
  • CD: Critical Difference (diagram)
  • FFG: Fusiform gyrus
  • IPL: Inferior Parietal Lobe
  • ITG: Inferior Temporal Lobe
  • MTG: Middle Temporal Gyrus